OpenStudy (anonymous):

ok,6 years ago,june was twice the age of sue and 8 years older than bill.if bill is now X years old.then in terms of X,what is sues current age?

OpenStudy (oblivious):

Writing out the equations described in this word problem: (J-6)=2*(S-6) (J-6)=(X-6)+8 X is Bill's age, J is june's, and S is sue's. We don't need Junes age so we set the two terms that are equal to J-6 equal to each other: 2*(S-6)=(X-6)+8 solve for S